OpenStudy (anonymous):

Center of mass of constant density p covering the region bounded below the paraola y=4-x^2 and above the x axis. I know i need to integrate to get the area but i'm not sure how i use it to get the center of mass

OpenStudy (anonymous):

use x1= 1/A\[\int\limits_{a}^{b}xf(x)dx\] and y1=\[1/A \int\limits_{a}^{b}1/2(f(x))^{2}dx\] Here, (x1,y1) is your center of mass...